5x-6y=-6
5x+51=y
5x-6(5x+51)=-6
5x-30x+306=-6
25x-300
x=-12,
and
y=5x+51
=-56+51
= my answer was -5 but, real answer is -9...
Can you please help me understand this?
Thank you so much!!!!

Found 2 solutions by rfer, PRMath:
Answer by rfer(16322)

I agree with everything you did down to x=-12
then I pluged -12 into one of the original equations.
5x-6y=-6
5*-12-6y=-6
-60-6y=-6
-6y=60-6
6y=-54
y=-9

Answer by PRMath(133)

5x-6y=-6
5x+51=y
5x-6(5x+51)=-6
5x-30x+306=-6
25x-300
x=-12,
Ok, all of the above is correct! :-) Good for you. Now you need to "plug" the -12 into your equation and that's where you made your mistake......
you had......

y=5x+51 <------this is correct.....

=-56+51 <------this is not correct. It should be:
y = 5(-12) + 51 (See where I "plugged" in the -12 for the "x" variable?
SOOOO now you solve:
y= 5(-12) +51
y= -60 + 51 (5 times -12 equals -60)
y= - 9 (-60 + 51 gives you -9)

Did that help? I hope so!
:-)